[Brahms-l] Shift report 20040202 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Mon Feb 02 2004 - 16:08:51 EST
------------------------------------------------------------
Date: Mon Feb  2 16:08:49 2004, Shift: 16:00-24:00
Supervisor: Ionut ARSENE, with: 
Log book#15 pages 48 - 48

Shift Summary:
------------------------------------------------------------
1. Usefull beam for 3 hours. In the rest of time we 
had problems with D1 and D2.

2. Runs: 10029 (374k trig 6 & 200k trig.3) with FS@3deg(1/1 field) and MRS@40 deg (B2255)
 Next runs are taken without FS data. All FS magnets are off.
 10033 (160k trig3) with MRS@40deg (A2255).
 10034 still running
   Pedestals:
 10030 - TPC pedestals
 10031 - non-TPC pedestal

3. No detector problems

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Feb 2 16:15:52 2004

This archive was generated by hypermail 2.1.8 : Mon Feb 02 2004 - 16:16:09 EST